OpenStudy (anonymous):

solve the eq.- cosA-sin3A=cos2A

OpenStudy (anonymous):

expand cos2A and sin3A

OpenStudy (dumbcow):

use trig identities to get everything in terms of cosA

OpenStudy (dumbcow):

cos(2A) = 2cos^2 - 1 , sin(2A) = 2sin(A)cos(A) sin(3A) = sin(2A+A) = sin(2A)cos(A) + sin(A)cos(2A)